what is the answer using the elimination/addition method to:
2w+3z=17
3w+4z=24

Answer by stanbon(75887)

answer using the elimination/addition method to:
2w+3z=17
3w+4z=24
----------
Multiply 1t by 3 and 2nd by 2 to get:
6w + 9z = 51
6w + 8z = 48
---------------
Subtract 2nd from 1st and solve for "z":
z = 3
---------
Substitute into 2w + 3z = 17 and solve for "w"
2w + 3*3 = 17
2w = 8
w = 4
---------
Ans: w = 4; z = 3
=========================
Cheers,
stan H.